# Contact Wasserstein Geodesics for Non-Conservative Schrödinger Bridges

### v2

@@ 4 The Non-Conservative Generalized Schrödinger Bridge
Guided NCGSB.    NCGSB (4) can be extended to the guided generation setting by introducing a guiding function $f$, which steers the generative process toward desired conditions at any chosen time (pmlr-v202-song23k; guo2024gradient). For a given time $t_{s}$, the guidance is expressed as $\color[rgb]{0.42,0.2,0.63}y=f(x^{t_{s}})$, with $x^{t_{s}}\sim\rho^{t_{s}}$. To enforce this form, the bridge $\rho^{t}$ is steered according to $\rho^{t}(x|y)=\frac{1}{Z}\,\rho^{t}(x)\,e^{\color[rgb]{0.42,0.2,0.63}-\|y-f(\bar{x}^{t_{s}})\|^{2}}$, where $Z$ is a normalization constant and $\bar{x}^{t_{s}}$ denotes a sample from the predicted guided distribution $\bar{\rho}^{t_{s}}$ conditioned on the current distribution $\rho^{t}$. By Bayes’ rule, the dynamics of the guided bridge $\rho^{t}(x|y)$ acquire an additional guidance term via the drift $v^{t}$, determined by $\color[rgb]{0.42,0.2,0.63}\|y-f(\bar{x}^{t_{s}})\|^{2}$. To perform a guided generation that enforces the constraint $y=f(x^{t_{s}})$ while preserving the underlying data manifold, we incorporate $y$ into the Lagrangian action constraint (4b) as (see App. C.3 for details),
$\partial_{t}\color[rgb]{0.06,0.3,0.7}z^{t}\color[rgb]{0.0,0.0,0.0}=\int_{\mathcal{M}}\left(\frac{1}{2}\|v^{t}(x)\|^{2}+U(x)+\color[rgb]{0.42,0.2,0.63}\|y-f(\bar{x}^{t_{s}})\|^{2}\color[rgb]{0.0,0.0,0.0}\right)\rho^{t}(x)\,dx\;-\;\gamma\color[rgb]{0.06,0.3,0.7}z^{t}\color[rgb]{0.0,0.0,0.0}.$ (5)
Wasserstein contact Hamiltonian flows and geodesics.    Analogously to mmGSB (2), understanding the dynamics of the optimality conditions in NCGSB (4) is essential for reformulating it as a geodesic computation. As detailed in App. C.2, we propose to leverage the contact Hamiltonian formalism (kholodenko2013applications), an extension of classical Hamiltonian mechanics to non-conservative systems (App. B.1), to model the dynamics of the NCGSB optimality conditions as Wasserstein contact Hamiltonian flows. This generalizes Prop. 1, since the contact Hamiltonian dynamics defines a geodesic but on the extended space $\mathcal{P}^{+}(\mathcal{M})\times\mathbb{R}$ (udricste2000geometric; testa2025geometric). The contact Hamiltonian optimality conditions are,
@@ 5 Contact Wasserstein Geodesics (CWG)
Approximate the evolution of the density $\rho^{t}\in\mathcal{P}^{+}(\mathcal{M})$ by a series of $K$ smooth parametrized pushforwards $T_{\theta^{k}}$, with $\theta^{k}$ belonging to a finite-dimensional space $\Theta$. If each pushforward $T_{\theta^{k}}$ is an immersion $T_{\theta^{k}}:\Theta\to\mathcal{T}\mathcal{P}^{+}(\mathcal{M})$, then the parameter space $\Theta$ can be endowed with a Riemannian structure via the pullback of the Wasserstein metric $g^{\mathcal{W}_{2}}$. Consequently, the contact Hamiltonian dynamics on $\mathcal{T}^{*}\mathcal{P}^{+}(\mathcal{M})\times\mathbb{R}$ can be represented in the reduced phase space $\Theta^{*}\times\mathbb{R}$, with the associated geodesic on $\mathcal{P}^{+}(\mathcal{M})\times\mathbb{R}$ projected onto $\Sigma\times\mathbb{R}$ (see App. D.1).
Proposition 2 allows us to transform the geodesic computation from the infinite-dimensional $\mathcal{P}^{+}(\mathcal{M})$ to a geodesic on a finite-dimensional parameterized space $\Sigma$ , such that the resulting geodesic flow on $\Sigma\times\mathbb{R}$ evolves under the pullback of the Jacobi metric $T_{\theta}^{*}\tilde{g}_{\mathrm{J}}=\Phi^{t_{k}}\,T_{\theta}^{*}g^{\mathcal{W}_{2}}$, where the scalar factor $\Phi^{t_{k}}=H(\rho^{t_{k}},S^{t_{k}},\color[rgb]{0.06,0.3,0.7}z^{t_{k}}\color[rgb]{0.0,0.0,0.0})-\mathcal{F}(\rho^{t_{k}})-\color[rgb]{.85,0.5,0.1}{\mathcal{B}(\rho^{t_{k}})}$, encodes the potential energy. Specifying the time evolution of $H^{t_{k}}$ determines a unique parameterized bridge on $\Sigma$. This formulation enables a tractable computation of geodesic flows to solve the NCGSB problem. Although different parameterizations $\{\theta^{k}\}_{k=0}^{K}$ may define distinct coordinate systems on $\Sigma$, the geodesics solutions remain equivalent and share the same energy (syrota2025identifying).
Geodesic computation.    The contact Wasserstein geodesic (CWG) corresponds to the discrete path $\{{\rho}_{\theta}^{t_{k}}\}_{k=0}^{K}$, that approximates the NCGSB solution. This is trained to reconstruct the available marginals while minimizing the geodesic energy under the pullback Jacobi metric $T_{\theta}^{*}\tilde{g}_{\mathrm{J}}=\Phi^{t_{k}}\,T_{\theta}^{*}g^{\mathcal{W}_{2}}$. The initial and final marginals, ${\rho}_{a}$ and ${\rho}_{b}$, are enforced at the path endpoints, corresponding to the ResNet outputs at times $t_{0}$ and $t_{K}$. All available intermediate marginals ${\rho}^{m}$ must appear at time points matching the ResNet discretization for the condition to be enforced.
CWG training happens in two stages (Alg. 1 in App. D.2): (1) we optimize the first ResNet block to match the initial marginal ${\rho}_{a}$, and (2) we find the optimal path by minimizing the loss,
@@
Here $d_{\mathcal{W}_{2}}$ denotes the Wasserstein-2 distance between probability distributions. In practice, this distance is approximated using empirical estimators based on samples drawn from the distributions. The complexity of this method is $\mathcal{O}\big(NK(T_{\text{sh}}+D(LW+\log N))\big)$, scaling linearly in dimension $D$ and nearly linearly in batch size $N$ (see App. D.4), rather than exponentially and quadratically (hong2025trajectory). Unlike chen2023deep; shen2025multimarginal, our CWG avoids costly iteration loops and is only weakly affected by the number of marginals.
@@
Guided contact Wasserstein geodesics.    In the conditional setting, the Lagrangian action dynamics (4b) in NCGSB (4) is augmented as in equation 5. Here, the scaling factor $\Phi^{t_{k}}=H(\rho^{t_{k}},S^{t_{k}},z^{t_{k}})-\mathcal{F}(\rho^{t_{k}})-\color[rgb]{.85,0.5,0.1}{\mathcal{B}(\rho^{t_{k}})}$ of the pullback Jacobi metric is augmented with the guidance term $\color[rgb]{0.42,0.2,0.63}\|y-f(\bar{x}^{t_{s}})\|^{2}$, to enforce the constraint $\color[rgb]{0.42,0.2,0.63}y=f(\bar{x}^{t_{s}})$ at time $t_{s}$ of the generative process. Under the ResNet parameterization, the desired distribution is approximated by $x^{t_{k_{s}}}\approx\bar{x}^{t_{s}}$ at time step $t_{k_{s}}$, and the Jacobi metric is modified as $\tilde{g}_{\mathrm{J}}^{\prime}=\big(\Phi^{t_{k}}+\color[rgb]{0.42,0.2,0.63}\color[rgb]{0.42,0.2,0.63}\|y-f(\bar{x}^{t_{s}})\|^{2}\color[rgb]{0.0,0.0,0.0}\big)\,g^{\mathcal{W}_{2}}$, with ${x}^{t_{k_{s}}}\sim{\rho}_{\theta}^{t_{k_{s}}}$. This penalizes geodesics crossing undesired regions at $t_{k_{s}}$. The loss for the guided optimization is
$\ell\!=\!d_{\mathcal{W}_{2}}^{2}(\rho_{\theta}^{t_{K}},\rho_{b})+\!\sum_{m=1}^{M}d_{\mathcal{W}_{2}}^{2}(\rho_{\theta}^{t_{k_{m}}},\rho_{m})+\!\sum_{k=1}^{K}\big(\Phi^{t_{k}}+\color[rgb]{0.42,0.2,0.63}\|y-f(\bar{x}^{t_{s}})\|^{2}\color[rgb]{0.0,0.0,0.0}\big)\,d_{\mathcal{W}_{2}}^{2}(\rho_{\theta}^{t_{k}},\rho_{\theta}^{t_{k-1}})+d_{\mathcal{W}_{2}}^{{}^{\prime}2}(\rho_{\theta}^{t_{t_{s}}},\rho_{s})$ (10)
@@
where the modified distance $d_{\mathcal{W}_{2}}^{{}^{\prime}2}$ measures deviations between the generated distribution ${\rho}_{\theta}^{t_{t_{s}}}$ and the intermediate marginal ${\rho}_{s}$ at $t_{s}$, while incorporating the penalty for samples $x_{s}$ that violate the guidance constraint $\color[rgb]{0.42,0.2,0.63}y=f(x_{s})$, c.f. App. E.1. In practice, this loss is optimized through a fine-tuning procedure applied to a model initially trained without any guidance.
@@
Proof of concept.    We demonstrate our framework on a 2D distribution-matching task and guided generation setting using the Two-Moons and Checkerboard benchmarks (flowsanddiffusions2025). These lack intermediate marginals $\{\rho_{m}\}_{m=1}^{M}$, and the initial distribution $\rho_{a}$ coincides with the reference distribution $\lambda$. Hence, only the second step of Alg. 1 is needed. Figure 4 shows that our method successfully generates the target distributions, and steers the generation to samples confined to a subset of the target space (here, the upper half). This guided behavior is achieved via the term $\color[rgb]{0.42,0.2,0.63}\|y-f(\bar{x}^{t_{s}})\|^{2}$, with $t_{s}\!=\!1$, $f$ measuring 2D sample positions, and $y$ defining the admissible region.
@@ 6 Results
We benchmark our approach against four established baselines summarized in Table 1. Further details of the experimental setups are provided in App. F.1.
@@
DSBM (shi2023diffusion)
GSBM (liu2024generalized)
SBIRR (shen2025multimarginal)
DM-SB (chen2023deep)
@@
First, we tackle a standard GSB task: computing a bridge evolving on a geometric manifold. We use the LiDAR scan of Mount Rainier (opentopo_rainier_2025) as the reference surface, and we aim to connect two marginals while remaining on the manifold and favoring low-altitude regions. These conditions are encoded into the potential function $U$ (see App. F.2). In this experiment, we do not model a physical system but instead compute an optimal transport map between the marginals under a conservative setting. Unlike our approach, DSBM and GSBM iteratively fit a deterministic path between the marginals, falling short on representing a posterior distribution. As a result, there is no guarantee that a Gaussian path remains on the manifold (Fig. 4). This leads to substantially higher-energy paths, as reflected by the Optimality metric in Table 6, and to inaccurate modeling of the target distribution, as indicated by the Feasibility metric (both metrics as defined in the baseline liu2024generalized). In contrast, our approach finds lower-energy solutions that respect the marginal constraints and converges significantly faster. Furthermore, our method uniquely supports guided generation, illustrated here by steering the probabilistic path to the right side of the mount (Fig. 4; see App. F.2 for quantitative results).
@@
Metric CWG (ours) GSBM DSBM
@@
tt (s) $\bm{1030_{\pm 50}}$ $73600_{\pm 3200}$ $19100_{\pm 900}$
Metric CWG (ours) GSBM DSBM
FID $\bm{{18.83}_{\pm 0.66}}$ ${40.23}_{\pm 1.95}$ ${149.78}_{\pm 0.81}$
tt (s) $\bm{{1090}_{\pm 40}}$ ${91100}_{\pm 8000}$ ${27400}_{\pm 2500}$
Single cell sequencing.    Next, we reconstruct stem cell differentiation dynamics from a series of isolated cellular snapshots. We use the Embryoid Body (EB) dataset from moon2019visualizing, which tracks cell state progression across five developmental stages $t_{0},t_{1},t_{2},t_{3},t_{4}]$. Cell differentiation is fundamentally a non-conservative biological process (zeevaert2020cell; kinney2014mesenchymal) and the ability to model energy-varying bridges is essential. To evaluate generalization in regions with no available data, we split the dataset into a training set $t_{0},t_{2},t_{4}]$ and a validation set $t_{1},t_{3}]$. Accordingly, the former contains the distributions $\{{\rho}_{a},{\rho}_{m_{2}},{\rho}_{b}\}$, while the latter contains $\{{\rho}_{m_{1}},{\rho}_{m_{3}}\}$. The geometry of the training distributions is encoded in the potential function $U$, which penalizes paths that stray from the observed data manifold. Minimizing $U$ ensures the learned bridge remains close to the data manifold, enabling effective generalization. The combination of the data manifold guidance and an energy-varying bridge allows our approach to outperform other mmSB baselines in both reconstruction accuracy and computation time. Quantitative results are reported in Table 6, with additional details and an ablation study on the importance of energy variation provided in App. F.3.
Figure 5: Predictions from CWG (ours, top), GSBM (middle), and DSBM (bottom). The red row shows marginal samples.
@@
Figure 6: FID scores at the validation time steps for three methods, evaluated over $9$ tests from 1981198120242024. Our CWG is significantly lower than the baselines.
@@
Image generation.    We also demonstrate our framework’s applicability to image generation tasks. Given sequences of images capturing the time evolution of physical phenomena, the model aims, from a single input image, to predict the most likely terminal state of the system, along with realistic intermediate frames at unobserved time steps. Specifically, we use the NOAA OISST v2 High Resolution Dataset (huang2021improvements), which provides daily sea surface temperature averages over multiple years, the BridgeData V2 (walke2023bridgedata) dataset, containing image snapshots of robotic manipulation tasks, and the Flickr-Faces-HQ (FFHQ) (karras2019style) dataset, an high-resolution ($1024\times 1024$) collection of human faces, where SB between distributions representing different categories (e.g., adults and children) enable unpaired image matching tasks. In the sea temperature prediction and robotic task reconstruction experiments, we compute the SB directly in the ambient image space. To ensure that the generated frames remain faithful to the underlying data distribution, we introduce a potential function $U$, that penalizes deviations from the learned data manifold. Specifically, for samples $x^{t}\sim\rho^{t}$, $U(x^{t})$ is defined as the reconstruction error, obtained via a VAE (song2025riemanniangeometric). In contrast, for the high-resolution unpaired image matching task, the SB is computed in a latent space, after encoding the images via the pre-trained ALAE encoder (pidhorskyi2020adversarial). Thus, no potential function is needed to align with the data distribution. Details on the energy behavior and extended results are provided in Apps. F.4, F.5, and F.6.
@@
For sea temperature prediction, we group data from 1981198120242024 into five-year intervals. Using heatmaps from January, May, and September (i.e., $t_{0},t_{2},t_{4}]$), our method predicts the temperature profiles of March and July (i.e., $t_{1},t_{3}]$). Our CWG produces cleaner, more accurate predictions than the baselines (Fig. 5). Since our framework operates efficiently in probability space and is not constrained by energy conservation, it achieves these results with an order of magnitude less computation (Table 6 for 20202020-$2024$; Fig. 6 for all years).

### v3

@@ 4 The Non-Conservative Generalized Schrödinger Bridge
Guided NCGSB.    NCGSB (4) can be extended to the guided generation setting by introducing a guiding function $f$, which steers the generative process toward desired conditions at any chosen time (Song et al., 2023; Guo et al., 2024). For a given time $t_{s}$, the guidance is expressed as $\color[rgb]{0.42,0.2,0.63}y=f(x^{t_{s}})$, with $x^{t_{s}}\sim\rho^{t_{s}}$. To enforce this form, the bridge $\rho^{t}$ is steered according to $\rho^{t}(x|y)=\frac{1}{Z}\,\rho^{t}(x)\,e^{\color[rgb]{0.42,0.2,0.63}-\|y-f({x}^{t_{s}})\|^{2}}$, where $Z$ is a normalization constant and ${x}^{t_{s}}$ denotes a sample from the predicted guided distribution ${\rho}^{t_{s}}(x|y)$. By Bayes’ rule, the dynamics of the guided bridge $\rho^{t}(x|y)$ acquire an additional guidance term via the drift $v^{t}$, determined by $\color[rgb]{0.42,0.2,0.63}\|y-f({x}^{t_{s}})\|^{2}$. To perform a guided generation that enforces the constraint $y=f(x^{t_{s}})$ while preserving the underlying data manifold, we incorporate $y$ into the Lagrangian action constraint (4b) as (see App. D.3 for details),
$\partial_{t}\color[rgb]{0.06,0.3,0.7}z^{t}\color[rgb]{0.0,0.0,0.0}=\int_{\mathcal{M}}\left(\frac{1}{2}\|v^{t}(x)\|^{2}+U(x)+\color[rgb]{0.42,0.2,0.63}\|y-f({x}^{t_{s}})\|^{2}\color[rgb]{0.0,0.0,0.0}\right)\rho^{t}(x)\,dx\;-\;\gamma\color[rgb]{0.06,0.3,0.7}z^{t}\color[rgb]{0.0,0.0,0.0}.$ (5)
Wasserstein contact Hamiltonian flows and geodesics.    Analogously to mmGSB (2), understanding the dynamics of the optimality conditions in NCGSB (4) is essential for reformulating it as a geodesic computation. As detailed in App. D.2, we propose to leverage the contact Hamiltonian formalism (Kholodenko, 2013), an extension of classical Hamiltonian mechanics to non-conservative systems (App. C.1), to model the dynamics of the NCGSB optimality conditions as Wasserstein contact Hamiltonian flows. This generalizes Prop. 1, since the contact Hamiltonian dynamics defines a geodesic but on the extended space $\mathcal{P}^{+}(\mathcal{M})\times\mathbb{R}$ (Udrişte, 2000; Testa et al., 2025). The contact Hamiltonian optimality conditions are,
@@ 5 Contact Wasserstein Geodesics (CWG)
Approximate the evolution of the density $\rho^{t}\in\mathcal{P}^{+}(\mathcal{M})$ by a series of $K$ smooth parametrized pushforwards $T_{\theta^{k}}$, with $\theta^{k}$ belonging to a finite-dimensional space $\Theta$. If each pushforward $T_{\theta^{k}}$ is an immersion $T_{\theta^{k}}:\Theta\to\mathcal{T}\mathcal{P}^{+}(\mathcal{M})$, then the parameter space $\Theta$ can be endowed with a Riemannian structure via the pullback of the Wasserstein metric $g^{\mathcal{W}_{2}}$. Consequently, the contact Hamiltonian dynamics on $\mathcal{T}^{*}\mathcal{P}^{+}(\mathcal{M})\times\mathbb{R}$ can be represented in the reduced phase space $\Theta^{*}\times\mathbb{R}$, with the associated geodesic on $\mathcal{P}^{+}(\mathcal{M})\times\mathbb{R}$ projected onto $\Sigma\times\mathbb{R}$ (see App. E.1).
Proposition 2 allows us to transform the geodesic computation from the infinite-dimensional $\mathcal{P}^{+}(\mathcal{M})$ to a geodesic on a finite-dimensional parameterized space $\Sigma$ , such that the resulting geodesic flow on $\Sigma\times\mathbb{R}$ evolves under the pullback of the Jacobi metric $T_{\theta}^{*}\tilde{g}_{\mathrm{J}}=\Phi^{t_{k}}\,T_{\theta}^{*}g^{\mathcal{W}_{2}}$, where the scalar factor $\Phi^{t_{k}}=H(\rho^{t_{k}},S^{t_{k}},\color[rgb]{0.06,0.3,0.7}z^{t_{k}}\color[rgb]{0.0,0.0,0.0})-\mathcal{F}(\rho^{t_{k}})-\color[rgb]{.85,0.5,0.1}{\mathcal{B}(\rho^{t_{k}})}$, encodes the potential energy. Specifying the time evolution of $H^{t_{k}}$ determines a unique parameterized bridge on $\Sigma$. This formulation enables a tractable computation of geodesic flows to solve the NCGSB problem. Although different parameterizations $\{\theta^{k}\}_{k=0}^{K}$ may define distinct coordinate systems on $\Sigma$, the geodesics solutions remain equivalent and share the same energy (Syrota et al., 2025).
Geodesic computation.    The contact Wasserstein geodesic (CWG) corresponds to the discrete path $\{{\rho}_{\theta}^{t_{k}}\}_{k=0}^{K}$, that approximates the NCGSB solution. This is trained to reconstruct the available marginals while minimizing the geodesic energy under the pullback Jacobi metric $T_{\theta}^{*}\tilde{g}_{\mathrm{J}}=\Phi^{t_{k}}\,T_{\theta}^{*}g^{\mathcal{W}_{2}}$. The initial and final marginals, ${\rho}_{a}$ and ${\rho}_{b}$, are enforced at the path endpoints, corresponding to the ResNet outputs at times $t_{0}$ and $t_{K}$. When available, intermediate marginals ${\rho}^{m}$ must appear at time points matching the ResNet discretization for the condition to be enforced.
CWG training happens in two stages (Alg. 1 in App. E.2): (1) we optimize the first ResNet block to match the initial marginal ${\rho}_{a}$, and (2) we find the optimal path by minimizing the loss,
@@
Here $d_{\mathcal{W}_{2}}$ denotes the Wasserstein-2 distance between probability distributions. In practice, this distance is approximated using empirical estimators based on samples drawn from the distributions. Convergence of the algorithm is guaranteed in App. E.3. Its complexity is $\mathcal{O}\big(NK(T_{\text{sh}}+D(LW+\log N))\big)$, scaling linearly in dimension $D$ and nearly linearly in batch size $N$ (see App. E.5), rather than exponentially or quadratically (Hong et al., 2025). Unlike Chen et al. (2023); Shen et al. (2025), our CWG avoids costly iteration loops and is only weakly affected by the number of marginals.
@@
Guided contact Wasserstein geodesics.    In the conditional setting, the Lagrangian action dynamics (4b) in NCGSB (4) is augmented as in equation 5. Here, the scaling factor $\Phi^{t_{k}}=H(\rho^{t_{k}},S^{t_{k}},z^{t_{k}})-\mathcal{F}(\rho^{t_{k}})-\color[rgb]{.85,0.5,0.1}{\mathcal{B}(\rho^{t_{k}})}$ of the pullback Jacobi metric is augmented with the guidance term $\color[rgb]{0.42,0.2,0.63}\|y-f({x}^{t_{s}})\|^{2}$, to enforce the constraint $\color[rgb]{0.42,0.2,0.63}y=f({x}^{t_{s}})$ at time $t_{s}$ of the generative process. Under the ResNet parameterization, the desired distribution is approximated by $x^{t_{k_{s}}}\approx{x}^{t_{s}}$ at time step $t_{k_{s}}$, and the Jacobi metric is modified as $\tilde{g}_{\mathrm{J}}^{\prime}=\big(\Phi^{t_{k}}+\color[rgb]{0.42,0.2,0.63}\color[rgb]{0.42,0.2,0.63}\|y-f({x}^{t_{s}})\|^{2}\color[rgb]{0.0,0.0,0.0}\big)\,g^{\mathcal{W}_{2}}$, with ${x}^{t_{k_{s}}}\sim{\rho}_{\theta}^{t_{k_{s}}}$. This penalizes geodesics crossing undesired regions at $t_{k_{s}}$. The loss for the guided optimization is
$\ell\!=\!d_{\mathcal{W}_{2}}^{2}(\rho_{\theta}^{t_{K}},\rho_{b})+\!\sum_{m=1}^{M}d_{\mathcal{W}_{2}}^{2}(\rho_{\theta}^{t_{k_{m}}},\rho_{m})+\!\sum_{k=1}^{K}\big(\Phi^{t_{k}}+\color[rgb]{0.42,0.2,0.63}\|y-f({x}^{t_{s}})\|^{2}\color[rgb]{0.0,0.0,0.0}\big)\,d_{\mathcal{W}_{2}}^{2}(\rho_{\theta}^{t_{k}},\rho_{\theta}^{t_{k-1}})+d_{\mathcal{W}_{2}}^{{}^{\prime}2}(\rho_{\theta}^{t_{t_{s}}},\rho_{s})$ (10)
@@
where the modified distance $d_{\mathcal{W}_{2}}^{{}^{\prime}2}$ measures deviations between the generated distribution ${\rho}_{\theta}^{t_{t_{s}}}$ and the intermediate marginal ${\rho}_{s}$ at $t_{s}$, while incorporating the penalty for samples $x_{s}$ that violate the guidance constraint $\color[rgb]{0.42,0.2,0.63}y=f(x_{s})$, c.f. App. F.1. In practice, this loss is optimized through a fine-tuning procedure applied to a model initially trained without any guidance.
@@
Proof of concept.    We demonstrate our framework on a 2D distribution-matching task and guided generation setting using the Two-Moons and Checkerboard benchmarks (Holderrieth and Erives, 2025). These lack intermediate marginals $\{\rho_{m}\}_{m=1}^{M}$, and the initial distribution $\rho_{a}$ coincides with the reference distribution $\lambda$. Hence, only the second step of Alg. 1 is needed. Figure 4 shows that our method successfully generates the target distributions, and steers the generation to samples confined to a subset of the target space (here, the upper half). This guided behavior is achieved via the term $\color[rgb]{0.42,0.2,0.63}\|y-f({x}^{t_{s}})\|^{2}$, with $t_{s}\!=\!1$, $f$ measuring 2D sample positions, and $y$ defining the admissible region.
@@ 6 Results
We benchmark our approach against four established baselines summarized in Table 1. Further details of the experimental setups are provided in App. G.1.
@@
DSBM (Shi et al., 2023)
SB-Flow (Bortoli et al., 2024)
GSBM (Liu et al., 2024)
SBIRR (Shen et al., 2025)
DM-SB (Chen et al., 2023)
@@
First, we tackle a standard GSB task: computing a bridge evolving on a geometric manifold. We use the LiDAR scan of Mount Rainier (OpenTopography, 2025) as the reference surface, and we aim to connect two marginals while remaining on the manifold and favoring low-altitude regions. These conditions are encoded into the potential function $U$ (see App. G.2). In this experiment, we do not model a physical system but instead compute an optimal transport map between the marginals under a conservative setting. Unlike our approach, DSBM and GSBM iteratively fit a deterministic path between the marginals, falling short on representing a posterior distribution. As a result, there is no guarantee that a Gaussian path remains on the manifold (Fig. 4). This leads to substantially higher-energy paths, as reflected by the Optimality metric in Table 3, and to inaccurate modeling of the target distribution, as indicated by the Feasibility metric (both metrics as defined in the baseline Liu et al. (2024)). In contrast, our approach finds lower-energy solutions that respect the marginal constraints and converges significantly faster. Furthermore, our method uniquely supports guided generation, illustrated here by steering the probabilistic path to the right side of the mount (Fig. 4; see App. G.2 for quantitative results).
@@
Single cell sequencing.    Next, we reconstruct stem cell differentiation dynamics from a series of isolated cellular snapshots. We use the Embryoid Body (EB) dataset from Moon et al. (2019), which tracks cell state progression across five developmental stages $t_{0},t_{1},t_{2},t_{3},t_{4}]$. Cell differentiation is fundamentally a non-conservative biological process (Zeevaert et al., 2020; Kinney et al., 2014) and the ability to model energy-varying bridges is essential. To evaluate generalization in regions with no available data, we split the dataset into a training set $t_{0},t_{2},t_{4}]$ and a validation set $t_{1},t_{3}]$. Accordingly, the former contains the distributions $\{{\rho}_{a},{\rho}_{m_{2}},{\rho}_{b}\}$, while the latter contains $\{{\rho}_{m_{1}},{\rho}_{m_{3}}\}$. The geometry of the training distributions is encoded in the potential function $U$, which penalizes paths that stray from the observed data manifold. Minimizing $U$ ensures the learned bridge remains close to the data manifold, enabling effective generalization. The combination of the data manifold guidance and an energy-varying bridge allows our approach to outperform other mmSB baselines in both reconstruction accuracy and computation time. Quantitative results are reported in Table 3, with additional details and an ablation study on the importance of energy variation provided in App. G.3.
Figure 5: Predictions generated by CWG (ours, top), GSBM (second), DSBM (third), and SB-Flow (fourth). The red row displays the corresponding reference samples.
@@
Metric CWG (ours) GSBM DSBM SB-Flow
@@
tt (60s) $\bm{17_{\pm 1}}$ $1227_{\pm 54}$ $318_{\pm 15}$ $83_{\pm 5}$
@@
Figure 6: FID scores at validation time steps for four methods, evaluated over 99 tests (1981–2024). Our CWG scores are significantly lower than baselines.
@@
Image generation.    We also demonstrate our framework’s applicability to image generation tasks. Given sequences of images capturing the time evolution of physical phenomena, the model aims, from a single input image, to predict the most likely terminal state of the system, along with realistic intermediate frames at unseen time steps. Specifically, we use the NOAA OISST v2 High Resolution Dataset (Huang et al., 2021), which provides daily sea surface temperature averages over multiple years, and the BridgeData V2 (Walke et al., 2023) dataset, containing image snapshots of robotic manipulation tasks. Our NCGSB problem can also be applied to unpaired image-matching tasks using general image datasets that do not describe dynamical systems, such as MNIST, EMNIST , and Flickr-Faces-HQ (FFHQ) (Karras et al., 2019). For the high-resolution unpaired image-matching task on the FFHQ dataset, the SB is computed in a latent space after encoding the images with the pre-trained ALAE encoder (Pidhorskyi et al., 2020). For all the other datasets, we compute the SB directly in the ambient image space. To ensure that the generated frames remain faithful to the underlying data distribution, we introduce a potential function $U$, that penalizes deviations from the learned data manifold. Following the baseline (Liu et al., 2024), for samples $x^{t}\sim\rho^{t}$, $U(x^{t})$ is defined as the reconstruction error, obtained via a VAE (Song and Itti, 2025). Details on the energy behavior and extended results are provided in Apps. G.4, G.5, G.6, and G.7.
@@
For the sea temperature prediction task, we cluster data from 1981198120242024 into five-year intervals. Using heatmaps from January, May, and September (i.e., $t_{0},t_{2},t_{4}]$), our method predicts the temperature profiles of March and July (i.e., $t_{1},t_{3}]$). Our CWG produces cleaner, more accurate predictions than the baselines (Fig. 6). Since our framework operates efficiently in probability space and is not constrained by energy conservation, it achieves these results with an order of magnitude less computation (Table 4 for 20202020-$2024$; Fig. 6 for all years).
@@
Metric CWG(ours) GSBM DSBM SB-Flow
FID $\bm{{19}_{\pm 1}}$ ${40}_{\pm 2}$ ${150}_{\pm 1}$ ${73}_{\pm 1}$
tt (h) $\bm{{0.5}_{\pm 0}}$ ${25.3}_{\pm 2.2}$ ${7.6}_{\pm 0.7}$ ${1.4}_{\pm 0.1}$
Table 5: FID score ($\downarrow$) and training time (tt) ($\downarrow$) in the robotic task reconstruction.